# CoFiI2P: Coarse-to-Fine Correspondences-Based Image-to-Point Cloud Registration

### v3

@@ IV-D1 Analyze of the I2P transformer
I2P transformer [37] with self-attention module and cross-attention module is crucial to image-to-point cloud alignment at the global level. In this part, we conduct ablation studies to assess the effectiveness of the I2P transformer. We train the CoFiI2P without any attention module as baseline. Then, the self-attention modules and cross-attention modules are added on the coarse level respectively. As shown in Table II, the baseline method performance drops significantly without any attention module. Besides, the self-attention module reduces the RRE about $1.06^{\circ}$ and the RTE about $0.47m$, and the cross-attention module reduces the RRE about $0.92^{\circ}$ and the RTE about $0.38m$ respectively. With both the self-attention and cross-attention modules, the CoFiI2P achieves the smallest registration error. Moreover, with both self-attention and cross-attention modules, the variance reduces by a large margin, which indicates that the I2P transformer block enhances both the accuracy and robustness.
@@
$2.88\pm 5.32$ $0.90\pm 2.44$
$1.82\pm 2.71$ $0.43\pm 0.75$
$1.96\pm 1.95$ $0.52\pm 0.59$
$\mathbf{1.28\pm 0.94}$ $\mathbf{0.33\pm 0.23}$
IV-D2 Analyze of the coarse-to-fine matching scheme
our CoFiI2P proposes to estimate coarse correspondences at super-point/-pixel level first and then generate dense correspondences at point-pixel level sequentially. To demonstrate that the progressive two-stage registration operates better than the one-stage registration used in previous I2P approaches, we conduct ablation experiments on the coarse-to-fine matching scheme. This ablation study employs the backbone with full I2P transformer blocks as baseline and evaluates the registration accuracy with only the coarse matching scheme or fine matching scheme. For coarse matching only, the matching pairs are established on the coarse level and remapped to the original resolution for pose estimation. By contrast, for fine matching only, matching pairs are established on the fine level directly, without guidance of coarse level correspondences. Experiment results in Table III show that removing either coarse matching stage or fine matching stage leads to higher registration error and variance. We indicate that coarse-level registration provides robust correspondences and fine-level registration provides accurate matching pairs. Combining the coarse-level and fine-level registration sequentially makes it easier to access the global optimal solution in I2P registration.
@@ IV-D2 Analyze of the coarse-to-fine matching scheme
$1.58\pm 1.64$ $0.41\pm 0.39$
$1.62\pm 2.89$ $0.44\pm 0.82$
$\mathbf{1.28\pm 0.94}$ $\mathbf{0.33\pm 0.23}$
IV-D3 Analysis of the point cloud density
given the significant impact of point cloud density on representative learning, we conducted ablation studies to examine this influence. The results in Table IV present registration accuracy and computational complexity for various point cloud densities. All experimental settings and hyperparameters remained consistent with those outlined in Section IV-A. The findings in Table IV reveal that as point cloud density decreases, the qualitative metrics deteriorate, while higher point cloud densities correspond to a sharp increase in computational complexity. It’s an intuitive observation that low-density point clouds lose local structured information, and high-density point clouds place a heavy computational burden. As a result, we opted for a compromise and selected 20480 points, striking a balance between efficiency and accuracy.
@@ IV-D3 Analysis of the point cloud density
5120 $10.68\pm 31.69$ $3.42\pm 22.85$ 37.42G
10240 $1.67\pm 2.07$ $0.42\pm 0.41$ 56.00G
20480 $1.28\pm 0.94$ $0.33\pm 0.23$ 93.80G
40960 $1.23\pm 1.30$ $0.32\pm 0.28$ 171.91G
@@ V Conclusion
In this paper, we propose CoFiI2P, a novel network for I2P registration. The core is the coarse-to-fine matching strategy, which establishes robust correspondence on the global level first and learns the precise correspondences on the local level progressively. Furthermore, the I2P transformer with self-attention and cross-attention module is introduced to enhance the global-aware ability in homogeneous and heterogeneous data. Compared with one-stage dense prediction and matching approaches, CoFiI2P filters out a large number of false correspondences and holds a safe lead in all the metrics. Extensive experiments on the KITTI dataset have demonstrated that CoFiI2P owns accuracy, robustness, and efficiency in various environments. We hope that the released code of CoFiI2P could motivate the related societies.

### v4

@@ IV-C1 Analysis of the I2P Transformer
I2P transformer [40] with self-attention module and cross-attention module is crucial to image-to-point cloud alignment at the global level. In this part, we conduct ablation studies to assess the effectiveness of the I2P transformer. We train the CoFiI2P without any attention module as baseline. Then, the self-attention modules and cross-attention modules are added on the coarse level respectively. As shown in Table IV, the baseline method performance drops significantly without any attention module. Besides, the self-attention module reduces the RRE to $1.78^{\circ}$ and the RTE about $0.41m$, and the cross-attention module reduces the RRE to $1.74^{\circ}$ and the RTE to $0.43m$ respectively. With both the self-attention and cross-attention modules, the CoFiI2P achieves the smallest registration error. Moreover, with both self-attention and cross-attention modules, the variance reduces by a large margin, which indicates that the I2P transformer block enhances both the accuracy and robustness.
IV-C2 Analysis of the Coarse-to-fine Matching Scheme
our CoFiI2P proposes to estimate coarse correspondences at super-point/-pixel level first and then generate fine correspondences at point-pixel level sequentially. We conduct ablation experiments on the coarse-to-fine matching scheme to demonstrate that the progressive two-stage registration operates better than the one-stage registration used in previous I2P approaches. This ablation study employs the backbone with full I2P transformer blocks as baseline and evaluates the registration accuracy with only coarse or fine matching schemes. For coarse matching only, matching pairs are established on the coarse level and remapped to the fine resolution for pose estimation. By contrast, for fine matching only, matching pairs are established on the fine level directly, without guidance of coarse level correspondences. Experiment results in Table IV show that removing either coarse matching stage or fine matching stage leads to higher registration error and variance. We indicate that coarse-level registration provides robust correspondences and fine-level registration provides accurate matching pairs. Combining the coarse-level and fine-level registration sequentially makes it easier to access the global optimal solution in I2P registration.
IV-C3 Analysis of the Point Cloud Density
Given the significant impact of point cloud density on representative learning, we conducted ablation studies to examine this influence. The results in Table IV present registration accuracy and computational complexity for various point cloud densities. The findings in Table IV reveal that as point cloud density decreases, the qualitative metrics deteriorate, while higher point cloud densities correspond to a sharp increase in computational complexity. It’s an intuitive observation that low-density point clouds lose local structured information, and high-density point clouds carry a heavy computational burden. As a result, we opted for a compromise and selected 20480 points, striking a balance between efficiency and accuracy.
@@ IV-C3 Analysis of the Point Cloud Density
$2.65\pm 4.79$ $0.87\pm 2.23$
${\color[rgb]{0,0,0}1.78}\pm{\color[rgb]{0,0,0}2.28}$ ${\color[rgb]{0,0,0}0.41}\pm{\color[rgb]{0,0,0}0.79}$
${\color[rgb]{0,0,0}1.74}\pm{\color[rgb]{0,0,0}1.50}$ ${\color[rgb]{0,0,0}0.43}\pm{\color[rgb]{0,0,0}0.35}$
$\mathbf{{\color[rgb]{0,0,0}1.14}\pm{\color[rgb]{0,0,0}0.78}}$ $\mathbf{{\color[rgb]{0,0,0}0.29}\pm{\color[rgb]{0,0,0}0.19}}$
@@
${\color[rgb]{0,0,0}1.35}\pm{\color[rgb]{0,0,0}1.17}$ ${\color[rgb]{0,0,0}0.34}\pm{\color[rgb]{0,0,0}0.26}$
${\color[rgb]{0,0,0}1.47}\pm{\color[rgb]{0,0,0}1.75}$ ${\color[rgb]{0,0,0}0.41}\pm{\color[rgb]{0,0,0}0.63}$
$\mathbf{{\color[rgb]{0,0,0}1.14}\pm{\color[rgb]{0,0,0}0.78}}$ $\mathbf{{\color[rgb]{0,0,0}0.29}\pm{\color[rgb]{0,0,0}0.19}}$
@@
5120 ${\color[rgb]{0,0,0}2.75}\pm{\color[rgb]{0,0,0}4.28}$ ${\color[rgb]{0,0,0}0.69}\pm{\color[rgb]{0,0,0}1.60}$ 37.42G
10240 ${\color[rgb]{0,0,0}1.69}\pm{\color[rgb]{0,0,0}1.43}$ ${\color[rgb]{0,0,0}0.39}\pm{\color[rgb]{0,0,0}0.30}$ 56.00G
20480 ${\color[rgb]{0,0,0}1.14}\pm{\color[rgb]{0,0,0}0.78}$ ${\color[rgb]{0,0,0}0.29}\pm{\color[rgb]{0,0,0}0.19}$ 93.80G
40960 ${\color[rgb]{0,0,0}1.00}\pm{\color[rgb]{0,0,0}0.70}$ ${\color[rgb]{0,0,0}0.26}\pm{\color[rgb]{0,0,0}0.17}$ 171.91G
@@ V Conclusion
This letter introduces CoFiI2P, a novel network designed for image-to-point cloud (I2P) registration. The proposed coarse-to-fine matching strategy first establishes robust global correspondences and then progressively refines precise local correspondences. Furthermore, the I2P transformer with self- and cross-attention modules is introduced to enhance the global-aware ability in homogeneous and heterogeneous data. Compared with existing one-stage dense prediction and matching approaches, CoFiI2P filters out a large number of false correspondences. Extensive experiments on the KITTI Odometry [36] and Nuscenes [37] dataset have demonstrated the superior accuracy, efficiency, and robustness of CoFiI2P in various environments. We hope the open-sourced CoFiI2P could benefit the relevant communities. In the near future, we will extend CoFiI2P to the unsupervised I2P registration.